OpenStudy (anonymous):

Write the equation of the line MN that contains Q(7,-2) and is parallel to y=2/5x-8/5

OpenStudy (anonymous):

I will medal and fan anyone please help!

OpenStudy (goldphenoix):

Alright. In this case, you will have to use the point slope formula, which is \[y -y_{2}=m(x-x _{2})\] Plug the coordinate in to get \[y-(-2)=m(x-(7))\] Now, we just need the slope. Well, y = 2/5 x - 8/5 has a slope of 2/5. FINALLY, we need to make the line parallel. To make two lines parallel, the lines have to have the same slope. Therefore, m = 2/5 \[y+2 = \frac{ 2 }{ 5 }(x-7)\] Clear? :)

OpenStudy (anonymous):

OH my gosh thank u so much!

OpenStudy (goldphenoix):

No problem.